A sample survey interviews an SRS of 207 college women. Suppose that 68% of all college women have been on a diet within the last 12 months. What is the probability that 73% or more of the women in the sample have been on a diet?

Use 4 decimal places.
Answer #1

Number of women on diet

Mean = n*p = 207*0.68 = 140.76

Standard dev = (n*p*q)^0.5 = (207*0.68*0.32)^0.5 = 6.711

Now 73% = 207*0.73 = 151.11

P(x>151.11) = 0.0618